At What temperature will a solution freeze if 200 grams of calcium chloride, CaCl2, is dissolved in 850 mL of water? kf H2O = 1.86°C/m


1
Expert's answer
2021-10-14T04:04:04-0400

∆T = KF·m

Molar mass of CaCl2 = 110.98 g/mol

200/110.98= 1.8mol

850/1000= 0.85kg

= 1.86× 0.85× 1.8= 2.8°C